A sample of gas has a pressure of 4.5 atm and a temperature of 325 K. What will be a new pressure if the temperature is
raised to 500K?
Law being used: ___________ FORMULA WORK ANSWER
Given 1:__________
Given 2:__________
Given 3:__________
1
Expert's answer
2020-04-18T07:00:32-0400

Solution:

Volume is constant


Law being used: Gay-Lussac’s Law


Given 1: __ P1 = 4.5 atm __

Given 2: __ T1 = 325 K __

Given 3: __T2 = 500 K __


FORMULA:

P1 / T1 = P2 / T2


WORK:

P1*T2 = P2*T1

(4.5 atm) * (500 K) = (P2) * (325 K)

P2 = (4.5 * 500) / 325 = 6.923 atm

P2 = 6.923 atm = 6.9 atm


ANSWER: P2 = 6.9 atm.